# Neutron star evolution by combining discontinuous Galerkin and finite volume methods

### v1

@@ Abstract
We present here a new hybrid scheme that combines a discontinuous Galerkin (DG) method with finite volume (FV) and finite difference (FD) methods. The computational mesh is divided into smaller elements that touch but do not overlap. Like a pure DG method, our new hybrid scheme requires information exchange only at the surface of neighboring elements. This avoids the need for ghostzones that are usually many points deep in traditional FV implementations. Furthermore, unlike traditional FV implementations, that require information exchange between each element and its 26 surrounding neighbors on non-cuboid meshes, our new hybrid method exchanges information only between each element and its six nearest neighbors. Through this reduction in communication, we aim to retain the high scalability of DG when using large supercomputers. The goal is to use DG in elements with smooth matter fields and to fall back onto the more robust FV/FD method in elements that contain non-smooth shocks or star surfaces. For this we devise trouble criteria to decide whether an element should be evolved with DG or FV/FD. We use the Nmesh program to implement and test the new scheme. We successfully evolve various single neutron star cases. These include the challenging cases of a neutron star initially in an unstable equilibrium migrating to a stable configuration and a boosted neutron star. These cases are simulated for the first time here in full 3D with general relativistic hydrodynamics using DG methods. We also describe additional numerical methods, such as the limiters and the atmosphere treatment we need for our simulations.
@@ I Introduction
The first gravitational wave (GW) signal observation in 2015 heralded the hundredth year commemoration of Einstein’s formulation of the theory of general relativity [1] and the associated equations of gravity [2]. This historic observation of a binary black hole (BBH) merger [3, 4] by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and the Virgo interferometer opened a new avenue of studying the universe through the GW spectrum. While other BBH merger observations [5, 6] followed, the first binary neutron star (BNS) merger was observed [7] only in 2017. This event, designated GW170817 by the LIGO collaboration, was also observed across the electromagnetic (EM) spectrum [8, 9, 10]. This observation of an EM counterpart for the GW signal launched the current epoch of multi-messenger astrophysics. The possibility of observing BNS mergers across multiple spectra make them especially interesting as GW sources. EM counterparts from BBH mergers can only be observed if the black holes are embedded in a dense matter environment that gets charged and heated during the inspiral and merger [11, 12, 13, 14], and hence are much rarer. BNS mergers and black hole-neutron star (BHNS) mergers [15] are thus much more promising and lucrative candidates for multi-messenger astrophysics. Furthermore, studying the neutron star (NS) related events also allows us to probe the related properties and phenomena such as equation of state (EoS) of matter at supranuclear densities [16, 17, 18], production of heavy elements via rapid neutron capture (r-process) nucleosynthesis [19, 20, 21], and value of the Hubble constant [22, 23].
@@
The sensitivity of the GW detectors has continuously increased [24, 25, 26, 27]. In addition, the LIGO-Virgo-KAGRA [28, 29, 30] detectors are projected to further increase in sensitivity [31]. More importantly, the sensitivity of the upcoming next generation [32] of GW detectors, namely Cosmic Explorer [33, 34], the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) [35], Einstein Telescope [36], LIGO Voyager [37], the Laser Interferometer Space Antenna (LISA) [38], NEMO [39], and TianQin [40] is expected to be yet much higher. We thus can look forward to GW observations with even higher signal-noise-ratio.
@@
GW signals observed are analyzed using template banks [41, 42] which in turn use results from databases of numerical simulations of CBCs, such as the CoRe [43], SXS [44], and RIT [45] databases. While older generation programs such as BAM [46, 47, 48], Einstein Toolkit [49, 50], SACRA-MPI [51], NRPy+ [52], and SpEC [53, 44] provide reliable simulation results, they cannot easily meet the demands for accuracy and speed for the upcoming GW detectors [54, 55, 56, 57, 58, 59]. In some cases this happens because a program may lack important features such as well parallelized and efficient adaptive mesh refinement (AMR). Yet in most cases, it is due to large overheads in the numerical methods used in older generation programs. The usual practice is to decompose the computational mesh into smaller domains or elements and to distribute these elements among available compute cores. However, since the physical system spans the entire mesh, the elements still need to communicate information between them. The large overhead comes from the need to exchange information many grid points deep, between nearest neighbor and sometimes even next to nearest neighbor elements, when one uses traditional finite volume (FV) or finite differencing (FD) methods. Additionally, for FV and FD schemes, achieving higher order error convergence is challenging, necessitating high resolution simulations, which in turn increase the cost again. Hence, fully utilizing the computing capability of modern high performance computers (HPCs), that contain hundreds of thousands or even million of compute cores, becomes difficult for these traditional methods.
@@
These challenges have motivated the creation of a new generation of NR programs, some of which we list here. Discontinuous Galerkin methods are used in bamps [60, 61], ExaHyPE [62, 63], and SpECTRE [64, 65]. Others like CarpetX [66], AsterX [67], GRaM-X [68], GR-Athena++ [69], AthenaK [70], Parthenon [71], GRChombo [72, 73], and Dendro-GR [74], explore leveraging graphics processing units (GPUs) while using FV or FD methods. Dendro-GR additionally uses wavelet-based adaptive mesh refinement. SPHINCS_BSSN [75] uses smooth particle hydrodynamics for matter evolution and FD for spacetime evolution.
@@
Nmesh [76] aims to be such a next-generation NR program. It uses a discontinuous Galerkin (DG) method to achieve its goal of high scalability and accuracy when running on modern HPCs. Introduced first by Reed and Hill [77] in 1973, the DG method took its rigorous form in a series of breakthrough articles by Cockburn and Shu [78, 79, 80, 81, 82]. Although historically employed in engineering and astrophysics, recent years have seen a growing interest in DG in the NR community [83, 60, 84, 61, 64, 85, 62, 63, 86, 87, 65]. DG offers two advantages: high order convergence when the fields are smooth and less communication overhead between neighboring elements distributed among different compute cores. The first comes from its spectral nature and the second comes from the nature of its boundary conditions between neighboring elements.
@@
However, simulation of neutron stars with full general relativistic hydrodynamics (GRHD) contain shocks and non-smooth features in the matter fields, in addition to smooth regions. While DG can handle shocks and non-smooth features between mesh element boundaries, it fails to adequately resolve them when these lie in the interior of an element. Such cases result in spurious Gibbs oscillations [88, 89, 90, 91] that drastically reduce accuracy. Additionally, DG is still subject to Godunov’s theorem [92] when dealing with discontinuities and thus has low order convergence when they are present. Hence, supplemental methods for handling these non-smooth features become necessary. While many non-linear limiters [93, 79, 82, 94, 95, 96] have been proposed for this purpose, they fall short [87, 65, 97] compared to the more robust traditional FV methods, with e.g. weighted essentially non-oscillatory (WENO) [98, 99] reconstruction, that are used for high resolution shock capturing (HRSC). This has motivated a different approach to the problem by using hybrid schemes that take advantage of DG’s higher accuracy at lower computational cost in regions where the matter fields are smooth, while falling back to the more robust FV+WENO formalism when handling non-smooth fields. Whether an element is evolved with DG or FV is determined by some sort of “trouble” detection criteria. The first implementation of this kind, to the best of our knowledge is presented in [100]. Subsequent examples of similar implementations can be found in [101, 102, 103, 104, 105, 63, 106, 87, 97]. A significant achievement in this approach, that has since been adopted by subsequent works, was the a posteriori approach in [103]. As in, trouble detection criteria are applied after a time step has been done. Then, if an element is troubled, and has been evolved with DG, the element is projected on to the FV or FD domain and the time step is redone. This is a significant difference from the limiters, which essentially tried to remove spurious Gibbs oscillations [88, 89, 90, 91] that inevitably arise when shocks are present. These, however, are better handled by FV methods with HRSC. Since the step is not redone to start from the physically sound values, the limiters approach inevitably retains some unphysciality that can grow over time, which is in contrast to the a posteriori approach.
@@
In this paper, we present a new FV/FD method that easily combines with the DG method. Our new scheme needs the same amount of communication as the DG method, and hence does not suffer from the communication overhead problem of traditional FV schemes described above. Part of our new FV/FD method is a new WENO reconstruction scheme, which we call “WENOZ2”, that uses a higher order WENOZ [107] method in the interior of an element and then drops the reconstruction order first to WENO3 [98] and then to even lower order as the element boundary is approached. We also devise certain trouble criteria (similar as in [87]) to detect elements with smooth matter fields, and use an a posteriori hybrid method that uses DG in smooth regions, but switches elements with non-smooth matter to use our new FV/FD method.
@@
We use dimensionless units where $G=c=M_{\odot}=1$. To convert to SI units, a dimensionless length has to be multiplied by $L_{0}=1476.6250$ m, a time by $T_{0}=4.9254909\times 10^{-6}$ s, a mass by $M_{\odot}=1.9884099\times 10^{30}$ kg, and a mass density by $6.1758285\times 10^{20}$ kg/m3. All quantities in this article are expressed in these units unless explicitly stated otherwise. For tensor notations, indices from Latin alphabet, such as $i$, run from $1$ to $3$ and denote spatial indices, while Greek indices, such as $\mu$, run from 00 to $3$ and denote spacetime indices.
II Equations of general relativistic hydrodynamics (GRHD)
@@ II.1 Metric evolution
Here $g_{\alpha\beta}$ is the spacetime metric and $\Gamma_{\alpha}=g^{\mu\nu}\Gamma_{\alpha\mu\nu}$ is the contracted Christoffel symbol. The equations are written in terms of the extra variables $\Pi_{\alpha\beta}:=-n^{\mu}{\partial}_{\mu}g_{\alpha\beta}$ and $\Phi_{i\alpha\beta}:={\partial}_{i}g_{\alpha\beta}$, that have been introduced to make the original second-order GHG system first-order in both time and space. Gauge conditions in the GHG system are specified by prescribing the gauge source function $H_{\alpha}$. We state our gauge choices for each of our simulation setups explicitly in Sec. IV. The lapse $\alpha$, shift $\beta^{i}$ and spatial metric $\gamma_{ij}$ come from the $3+1$ decomposition in Eq. (1).
The GHG evolution equations also contain extra terms for constraint damping that are multiplied with the factors $\gamma_{0}$, $\gamma_{1}$, and $\gamma_{2}$. For some of the runs, we set these factors to the constant values of $\gamma_{0}=0.1$, $\gamma_{1}=-1$, and $\gamma_{2}=1$. Yet in most other cases, we set them to [97]
@@ II.2 Evolution of hydrodynamic fields
The evolution equations for the matter are typically brought into conservative from
@@
Here $D$ is rest-mass density, $\tau$ the internal energy density, $S_{i}$ the momentum density as seen by Eulerian observers. The last two equations also contain the specific enthalpy given by
@@
To close the evolution system, we have to specify an equation of state (EoS) for the fluid, i.e. an equation of the form
@@ III Numerical methods
We use the same DG implementation that we use in [76]. However, we find that this DG method is not enough for simulating cases of boosted stars with relatively higher velocities or unstable stars that rapidly expand or contract. For matter fields with fast moving non-smooth features or even shocks, the much more robust FV method with e.g. WENO [98] is necessary. Regions containing non-smooth features in the matter fields are classified as “troubled”. In these troubled regions we will use an FV method for the matter fields and an FD method for the gravitational fields. We will refer to this combination as the “FV/FD” method. However, we still want to use DG for all fields in non-troubled regions to leverage its higher accuracy and lower computational cost. Hence, we opt for a hybrid DG+FV/FD method. This has been explored in other works before [101, 102, 103, 62, 87], but the specifics of combining the two schemes differ between those works and ours. Our FV/FD is more compatible with our DG implementation and also our goal of high scalability.
@@ III.1 The DG method
$=s_{q_{1}q_{2}q_{3}}$displaystyle-- $\frac{\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}_{q_{1}q_{2}q_{3}}}}{w_{q_{1}}}F_{q_{1}q_{2}q_{3}}(\delta_{q_{1}0}+\delta_{q_{1}N})$ (26)
@@
On the other hand, for a set of non-conservative equations of the form
@@
$=s_{q_{1}q_{2}q_{3}}$displaystyle-- $\frac{\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}_{q_{1}q_{2}q_{3}}}}{w_{q_{1}}}G_{q_{1}q_{2}q_{3}}(\delta_{q_{1}0}+\delta_{q_{1}N})$ (29)
@@
III.2 Modified finite volume (FV) / finite difference (FD) method
In order to deal with shocks or other non-smooth behavior such as star surfaces it is often better to use finite volume or finite difference methods. Here we explain how we can modify Eqs. (III.1) and (28) for such cases. First we will change our mesh and use $N+1$ equally spaced grid points with grid spacing $h=2/N$.

### v2

@@ Abstract
We present here a new hybrid scheme that combines a discontinuous Galerkin (DG) method with compact finite volume (FV) and finite difference (FD) methods. The computational mesh is divided into smaller elements that touch but do not overlap. Like a pure DG method, our new hybrid scheme requires information exchange only at the surface of neighboring elements. This avoids the need for ghost zones that are usually many points deep in traditional FV implementations. Furthermore, unlike traditional FV implementations, that require information exchange between each element and its 26 surrounding neighbors on noncuboid meshes, our new hybrid method exchanges information only between each element and its six nearest neighbors. Through this reduction in communication, we aim to retain the high scalability of DG when using large supercomputers. In addition, the information exchange between adjacent elements is much simpler than in a traditional FV implementation, because we always have grid points at the interface, so that only surface interpolation is required. As a result it is much easier to implement adaptive mesh refinement. The goal is to use DG in elements with smooth matter fields and to fall back onto the more robust FV/FD method in elements that contain nonsmooth shocks or star surfaces. For this we devise trouble criteria to decide whether an element should be evolved with DG or FV/FD. We use the Nmesh program to implement and test the new scheme. We successfully evolve various single neutron star cases. These include the challenging cases of a neutron star initially in an unstable equilibrium migrating to a stable configuration and a boosted neutron star. These cases are simulated for the first time here in full 3D with general relativistic hydrodynamics using DG methods. We also describe additional numerical methods, such as the limiters and the atmosphere treatment we need for our simulations.
@@ I Introduction
The first gravitational wave (GW) signal observation in 2015 heralded the hundredth year commemoration of Einstein’s formulation of the theory of general relativity [1] and the associated equations of gravity [2]. This historic observation of a binary black hole (BBH) merger [3, 4] by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and the Virgo interferometer opened a new avenue of studying the universe through the GW spectrum. While other BBH merger observations [5, 6] followed, the first binary neutron star (BNS) merger was observed [7] only in 2017. This event, designated GW170817 by LIGO collaboration, was also observed across the electromagnetic (EM) spectrum [8, 9, 10]. This observation of an EM counterpart for the GW signal launched the current epoch of multimessenger astrophysics. The possibility of observing BNS mergers across multiple spectra make them especially interesting as GW sources. EM counterparts from BBH mergers can only be observed if the black holes are embedded in a dense matter environment that gets charged and heated during the inspiral and merger [11, 12, 13, 14], and hence are much rarer. BNS mergers and black hole-neutron star mergers [15] are thus much more promising and lucrative candidates for multimessenger astrophysics. Furthermore, studying the neutron star (NS) related events also allows us to probe the related properties and phenomena such as equation of state (EoS) of matter at supranuclear densities [16, 17, 18], production of heavy elements via rapid neutron capture (r-process) nucleosynthesis [19, 20, 21], and value of the Hubble constant [22, 23].
@@
The sensitivity of the GW detectors has continuously increased [24, 25, 26, 27]. In addition, the LIGO-Virgo-KAGRA [28, 29, 30] detectors are projected to further increase in sensitivity [31]. More importantly, the sensitivity of the upcoming next generation [32] of GW detectors, namely Cosmic Explorer [33, 34], the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) [35], Einstein Telescope [36], LIGO Voyager [37], the Laser Interferometer Space Antenna (LISA) [38], NEMO [39], and TianQin [40] is expected to be yet much higher. We thus can look forward to GW observations with even higher signal-to-noise ratio.
@@
GW signals observed are analyzed using template banks [41, 42] which in turn use results from databases of numerical simulations of compact binary coalescences (CBCs), such as the CoRe [43], SXS [44], and RIT [45] databases. While older generation programs such as BAM [46, 47, 48], Einstein Toolkit [49, 50], SACRA-MPI [51], NRPy+ [52], and SpEC [53, 44] provide reliable simulation results, they cannot easily meet the demands for accuracy and speed for the upcoming GW detectors [54, 55, 56, 57, 58, 59]. In some cases this happens because a program may lack important features such as well-parallelized and efficient adaptive mesh refinement (AMR). Yet in most cases, it is due to large overheads in the numerical methods used in older generation programs. The usual practice is to decompose the computational mesh into smaller domains or elements and to distribute these elements among available compute cores. However, since the physical system spans the entire mesh, the elements still need to communicate information between them. The large overhead comes from the need to exchange information many grid points deep, between nearest neighbor and sometimes even next to nearest neighbor elements, when one uses traditional finite volume (FV) or finite differencing (FD) methods. Additionally, for FV and FD schemes, achieving higher-order error convergence is challenging, necessitating high-resolution simulations, which in turn increase the cost again. Hence, fully utilizing the computing capability of modern high-performance computers (HPCs), that contain hundreds of thousands or even million of compute cores, becomes difficult for these traditional methods.
@@
These challenges have motivated the creation of a new generation of numerical relativity (NR) programs, some of which we list here. Discontinuous Galerkin methods are used in bamps [60, 61], ExaHyPE [62, 63], and SpECTRE [64, 65]. Others like CarpetX [66], AsterX [67], GRaM-X [68], GR-Athena++ [69], AthenaK [70], Parthenon [71], GRChombo [72, 73], and Dendro-GR [74], explore leveraging graphics processing units while using FV or FD methods. Dendro-GR additionally uses wavelet-based adaptive mesh refinement. SPHINCS_BSSN [75] uses smooth particle hydrodynamics for matter evolution and FD for spacetime evolution.
@@
Nmesh [76] aims to be such a next-generation NR program. It uses a discontinuous Galerkin (DG) method to achieve its goal of high scalability and accuracy when running on modern HPCs. Introduced first by Reed and Hill [77] in 1973, the DG method took its rigorous form in a series of breakthrough articles by Cockburn and Shu [78, 79, 80, 81, 82]. Although historically employed in engineering and astrophysics, recent years have seen a growing interest in DG in the NR community [83, 60, 84, 61, 64, 85, 62, 63, 86, 87, 65]. DG offers two advantages: high-order convergence when the fields are smooth and less communication overhead between neighboring elements distributed among different compute cores. The first comes from its spectral nature and the second comes from the nature of its boundary conditions between neighboring elements.
@@
However, simulation of neutron stars with full general relativistic hydrodynamics (GRHD) contain shocks and nonsmooth features in the matter fields, in addition to smooth regions. While DG can handle shocks and nonsmooth features between mesh element boundaries, it fails to adequately resolve them when these lie in the interior of an element. Such cases result in spurious Gibbs oscillations [88, 89, 90, 91] that drastically reduce accuracy. Additionally, DG is still subject to Godunov’s theorem [92] when dealing with discontinuities and thus has low-order convergence when they are present. Hence, supplemental methods for handling these nonsmooth features become necessary. While many nonlinear limiters [93, 79, 82, 94, 95, 96] have been proposed for this purpose, they fall short [87, 65, 97] compared to the more robust traditional FV methods, with e.g. weighted essentially nonoscillatory (WENO) [98, 99] reconstruction, that are used for high-resolution shock capturing (HRSC). This has motivated a different approach to the problem by using hybrid schemes that take advantage of DG’s higher accuracy at lower computational cost in regions where the matter fields are smooth, while falling back to the more robust FV+WENO formalism when handling nonsmooth fields. Whether an element is evolved with DG or FV is determined by some sort of “trouble” detection criteria. The first implementation of this kind, to the best of our knowledge is presented in [100]. Subsequent examples of similar implementations can be found in [101, 102, 103, 104, 105, 63, 106, 87, 97]. A significant achievement in this approach, that has since been adopted by subsequent works, was the a posteriori approach in [103]. As in, trouble detection criteria are applied after a time step has been done. Then, if an element is troubled, and has been evolved with DG, the element is projected on to the FV or FD domain and the time step is redone. This is a significant difference from the limiters, which essentially tried to remove spurious Gibbs oscillations [88, 89, 90, 91] that inevitably arise when shocks are present. These, however, are better handled by FV methods with HRSC. Since the step is not redone to start from the physically sound values, the limiters approach inevitably retains some unphysciality that can grow over time, which is in contrast to the a posteriori approach.
@@
In this paper, we present a new FV/FD method that easily combines with the DG method. Our new scheme needs the same amount of communication as the DG method, and hence does not suffer from the communication overhead problem of traditional FV schemes described above. Part of our new FV/FD method is a new WENO reconstruction scheme, which we call “WENOZ2,” that uses a higher order WENOZ [107] method in the interior of an element and then drops the reconstruction order first to WENO3 [98] and then to even lower order as the element boundary is approached. We also devise certain trouble criteria (similar as in [87]) to detect elements with smooth matter fields, and use an a posteriori hybrid method that uses DG in smooth regions, but switches elements with nonsmooth matter to use our new FV/FD method.
@@
We use dimensionless units where $G=c=M_{\odot}=1$. To convert to Système international (SI) units, a dimensionless length has to be multiplied by $L_{0}=1476.6250$ m, a time by $T_{0}=4.9254909\times 10^{-6}$ s, a mass by $M_{\odot}=1.9884099\times 10^{30}$ kg, and a mass density by $6.1758285\times 10^{20}$ kg/m3. All quantities in this article are expressed in these units unless explicitly stated otherwise. For tensor notations, indices from Latin alphabet, such as $i$, run from $1$ to $3$ and denote spatial indices, while Greek indices, such as $\mu$, run from 0 to $3$ and denote spacetime indices.
II Equations of general relativistic hydrodynamics
@@ II.1 Metric evolution
Here $g_{\alpha\beta}$ is the spacetime metric and $\Gamma_{\alpha}=g^{\mu\nu}\Gamma_{\alpha\mu\nu}$ is the contracted Christoffel symbol. The equations are written in terms of the extra variables $\Pi_{\alpha\beta}:=-n^{\mu}{\partial}_{\mu}g_{\alpha\beta}$ and $\Phi_{i\alpha\beta}:={\partial}_{i}g_{\alpha\beta}$, that have been introduced to make the original second order GH system first order in both time and space. Gauge conditions in the GH system are specified by prescribing the gauge source function $H_{\alpha}$. We state our gauge choices for each of our simulation setups explicitly in Sec. IV. The lapse $\alpha$, shift $\beta^{i}$ and spatial metric $\gamma_{ij}$ come from the $3+1$ decomposition in Eq. (1).
The GH evolution equations also contain extra terms for constraint damping that are multiplied with the factors $\gamma_{0}$, $\gamma_{1}$, and $\gamma_{2}$. For some of the runs, we set these factors to the constant values of $\gamma_{0}=0.1$, $\gamma_{1}=-1$, and $\gamma_{2}=1$. Yet in most other cases, we set them to [97]
@@ II.2 Evolution of hydrodynamic fields
The evolution equations for the matter are typically brought into conservative form
@@
Here $D$ is rest-mass density, $\tau$ the internal energy density, and $S_{i}$ the momentum density as seen by Eulerian observers. The last two equations also contain the specific enthalpy given by
@@
To close the evolution system, we have to specify an EoS for the fluid, i.e. an equation of the form
@@ III Numerical methods
We use the same DG implementation that we use in [76]. However, we find that this DG method is not enough for simulating cases of boosted stars with relatively higher velocities or unstable stars that rapidly expand or contract. For matter fields with fast moving nonsmooth features or even shocks, the much more robust FV method with e.g. WENO [98] is necessary. Regions containing nonsmooth features in the matter fields are classified as “troubled”. In these troubled regions we will use an FV method for the matter fields and an FD method for the gravitational fields. We will refer to this combination as the “FV/FD” method. However, we still want to use DG for all fields in nontroubled regions to leverage its higher accuracy and lower computational cost. Hence, we opt for a hybrid DG+FV/FD method. This has been explored in other works before [101, 102, 103, 62, 87], but the specifics of combining the two schemes differ between those works and ours. Our FV/FD is more compatible with our DG implementation and also our goal of high scalability.
@@ III.1 The DG method
$=s_{q_{1}q_{2}q_{3}}$ - $\frac{\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}_{q_{1}q_{2}q_{3}}}}{w_{q_{1}}}F_{q_{1}q_{2}q_{3}}(\delta_{q_{1}0}+\delta_{q_{1}N})$ (26)
@@
On the other hand, for a set of nonconservative equations of the form
@@
$=s_{q_{1}q_{2}q_{3}}$ - $\frac{\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}_{q_{1}q_{2}q_{3}}}}{w_{q_{1}}}G_{q_{1}q_{2}q_{3}}(\delta_{q_{1}0}+\delta_{q_{1}N})$ (29)
@@
III.2 Modified finite volume / finite difference method
In order to deal with shocks or other nonsmooth behavior such as star surfaces it is often better to use finite volume or finite difference methods. Here we explain how we can modify Eqs. (III.1) and (III.1) for such cases. First we will change our mesh and use $N+1$ equally spaced grid points with grid spacing $h=2/N$.